OpenStudy (anonymous):

draw the line with the given slope and y-intercept m=2 b= -4

OpenStudy (user917):

we know y=ax+b

OpenStudy (user917):

a is the slope and b is y-intercept

OpenStudy (user917):

so what will be now the equation of the line?

OpenStudy (anonymous):

@user917 y=2x+(-4)

OpenStudy (user917):

Excellent! Its y=2x-4

OpenStudy (anonymous):

Oh okay I see. It's minus and not addition? Whats the next step?

OpenStudy (user917):

its addition (y=ax+b) but +(-4)=-4

OpenStudy (anonymous):

oh okay i see

OpenStudy (user917):

There is no next step, you drew the line, this is the answer = y=2x-4

OpenStudy (anonymous):

How would i show that on a graph?

OpenStudy (anonymous):

 like this ? @user917

OpenStudy (user917):

yes cool that is great! you take 2 arbitrary points x and you calculate its y.. in this case for x-0 y=-4 and for x=2 y=0, you put this two point on the graph like you did and then you draw a line passing through these 2 points. Great Job :)

OpenStudy (anonymous):

@user917 Sweet! Alright, thank you so much! :)